OpenStudy (anonymous):

The area of a triangle is equal to 48 cm2 and two of its dies measure 12 cm and 9.34, respectively. Find the measure of the included OBTUSE angle (in degrees) of the given sides. Round off your answer to two decimal places

OpenStudy (michele_laino):

is your angle enclose between those sides?

OpenStudy (michele_laino):



OpenStudy (anonymous):

yup

OpenStudy (anonymous):

i have an answer of 10.75 but not so sure of how to get the proper obtuse angle.

OpenStudy (anonymous):

and when I rechecked and went through my calculation I now have 121.07..

OpenStudy (michele_laino):

ok! the you have to apply this formula: \[A = \frac{1}{2}12 \cdot 9.34 \cdot \sin x\] where A is thearea of your triangle, namely A = 48 cm^2

OpenStudy (bradely):

use formula A =1/2(ab)sinC 48=1/2(12*9.34)sinC 56.04sinC =48 sinC = 0.86 Use calculator C=59 Source: http://www.mathskey.com/question2answer/

OpenStudy (michele_laino):



OpenStudy (anonymous):

thanks a lot just got a bit confused -.-

OpenStudy (michele_laino):

please where you are confused?

OpenStudy (michele_laino):

please what is \[\frac{1}{2}12 \cdot 9.34 = ...?\]

OpenStudy (anonymous):

it's 56.04 then divide 48 by 56.04 = 0.857 then arcsin(0.857) =58.98